# The Open Journal of Astrophysics Sulphur abundances in star-forming regions from optical emission lines: A new approach based on photoionization models consistent with the direct method

### v1

@@ 4.1 Comparison with the control sample
In Figure 3 we show the comparison between the abundances calculated by Díaz & Zamora (2022) with the results from HCm when all possible lines, including auroral lines, are used as input by the code. We also list in Table LABEL:disp the mean offsets and the standard deviations of these comparisons. As it can be seen, the agreement in both plots is quite good, without any neither discontinuity nor break over a wide range of metallicity, and only with some little deviations that can be attributed to several factors including the atomic coefficients, the assumed relations between the electron temperatures associated to each ion, or the assumed ICFs for each element. These factors are arbitrary and can lead to deviations that are usually lower than the observational uncertainties.
In absence of any auroral line, which is the most common situation in weak regions or for metal-rich objects, where the cooling is much more efficient and the emissivities of all CELs is lower, the code can also provide a solution with its corresponding uncertainty. However, additional assumptions must be taken in order to provide a solution consistent with the direct method. As discussed in Pérez-Montero (2014), this is done by assuming an empirical relation between O/H and to constrain the grids, in such a way that models with lower metallicity have in average higher excitation and, on the contrary, low excitation regions are compatible with metal-rich regions. This assumption helps breaking the degeneracy of certain emission-line ratios such as R23, with metallicity. A similar constrain is considered by the code when N/O cannot be calculated, so [N ii] emission-lines can only be used to derive O/H after assuming a certain arbitrary relation between O/H and N/O. As discussed in Pérez-Montero (2014) and several other works describing HCm, this assumption is usually made in all strong-line methods based on N ii emission lines in the optical, so it can lead to non-negligible mistakes if the studied object does not follow the foreseen adopted relation. Anyway, these relations can be edited or replaced in the code to account for different families of objects or prescriptions. For instance, as studied in Pérez-Montero et al. (2021), for EELGs, the assumed relation between O/H and is slightly different, pointing towards slightly higher values for lower metallicity.
@@
4.2 An assessment of the empirical relation between O/H and
@@
One of the basements of the above described methodology using the HCm code to derive chemical abundances when no nebular-to-auroral line ratios are provided in the input is the assumed empirical anticorrelation between O/H and , what is used by the code constraining the grid of models to avoid the double-valued relations of certain emission-line ratios with O/H (e.g. R23, Pagel et al. 1979). Thus, it is worth to make a deeper examination on the physical basis of this assumption.
@@
In left panel of Figure 5 we show the relation between the values of O/H and found by HCm for the objects in the control sample for which at least one auroral line is used as an input. In this case, the code does not constrain the grid of models used to perform the calculations and thus no previous assumption is made about the relation between O/H and in this case. . As it can be seen, the code finds the expected anticorrelation already found for a larger sample in Pérez-Montero (2014), in such a way that objects with a lower metallicity have in average higher values for log . This relation was also obtained using a larger sample in Pérez-Montero (2014) and is represented by the solid line shown in the same figure. This line also encompasses The set of models used by the code to perform the calculations when no auroral line is given as input.
@@
The subsample from Díaz & Zamora (2022) used in this work shows an anticorrelation between O/H and , as found by HCm, somehow less clear than in the case of the sample studied in Pérez-Montero (2014) ( = -0.48), maybe owing to the higher weight in the sample analyzed here of the objects at large metallicity, for which the trend does not appear. However, despite that some objects lie out of the limits designed to constrain the grid of models, the agreement between the abundances derived by the code when no auroral lines are given using a constrained grid is quite reasonable, as can be seen in Fig. 4 and inspecting the offsets and dispersions listed in Table LABEL:disp.
@@
However, focusing on this observed O/H - anticorrelation itself, there is not a consensus in the literature about if this correlation is real or has a physical basis. On one hand, several authors (e.g. Dopita & Evans 1986; Maier et al. 2006; Morisset et al. 2016; Garner et al. 2025) have reported the same anticorrelation in samples of H ii regions by fitting high-to-low excitation emission-line ratios with photoionization models. On the other hand, some other authors found a lack of correlation or even a positive correlation between O/H and U (e.g. Kennicutt & Garnett 1996; Dors et al. 2011; Kreckel et al. 2019; Zinchenko et al. 2019) and/or alert on the strong dependence of the results on the initial assumptions made in the photoionization models used to derive these properties (e.g. Ji & Yan 2022).
@@
In the several papers describing the different versions of HCm, it is stated that the derived U values are subject to the definition of the assumed geometry or the adopted input SED, so its final value is just a convention taken to compare heterogeneous samples or to fix an empirical relation between the metal content of the gas and the observed behaviour of emission-line ratios very dependent on the gas excitation, such as O2O3 or S2S3, as explained in the above sections. However, a grid of models exploring a variation of properties such as the shape of the incident SED, the geometry, or the assumed dust-to-gas mass ratio is required to correctly interpret from a physical basis the observed anticorrelation between metallicity and , as all these other properties have a non-negligible impact on the observed emission-line ratios.
@@
It has been demonstrated that making small modifications333As explored in Pérez-Montero (2014) variations of the age of the popstar Mollá et al. (2009) stellar atmospheres used in the models do not lead to deviations in the derived oxygen or nitrogen abundances larger than 0.02 dex. In addition, in Pérez-Montero et al. (2019a),, for AGN, a change in the assumed parameter shaping the incident SED, do not alter the derived abundances. Finally, in Pérez-Montero et al. (2021), for Extreme Emission Line Galaxies (EELGs), it is proven that changes in the properties of the assumed BPASS Eldridge et al. (2017) model atmospheres, such as age, IMF or upper mass limit, only imply variations in the derived log , but neither for N/O nor for O/H. on the assumed incident SED of the models does not imply changes beyond the reported errors for the results from HCm on the derived abundances, what is corroborated by the agreement found when these results are compared with those coming from the direct method, Nevertheless, a closer inspection of the real nature of the derived values in HCm can only be made by means of an alternative approach exploring the nature of the incident field of radiation. For instance, this can be performed by using the code HCm-Teff Pérez-Montero et al. (2019b), with a very similar methodology to HCm, but using as inputs the metal content of the gas, and at least two emission-line ratios to simultaneously fix the ionization parameter () and the effective equivalent temperature of the ionizing source (), thus providing an additional perspective to the observed anticorrelation between O/H and the obtained parameter when the SED is fixed. The study of the control sample studied in this work allows this study, as it has been selected to simultaneously have emission-line ratios, O2O3 and S2S3 involved in the softness diagram Pérez-Montero & Vílchez (2009), what allows HCm-Teff to extract both and .
@@
The anticorrelation between O/H and was already pointed by Pérez-Montero et al. (2019b) analyzing the H ii regions in NGC 628, M 51, and M 101, where the presence of a negative radial gradient of metallicity were obtained in coincidence with a positive radial gradient of the hardness of the incident field of radiation. This relation was also indirectly pointed out by other authors (e.g. Dopita et al. 2006; Eldridge et al. 2011), who state the the relation observed with when this is derived using emission-line ratios could be the consequence of a higher opacity in the atmospheres of the ionizing stars. At same time, other empirically found relations between metallicity and other properties of the ionizing stellar population have been found in other works, such as Initial Mass Function (IMF, Martín-Navarro et al. 2015), SFR or sSFR (e.g. Dopita et al. 2014; Mingozzi et al. 2020), or age Pellegrini et al. (2020), underlining the anticorrelation between metallicity and the hardness of the incident field of radiation, and not for the ionization parameter.
Nevertheless, a more precise and robust characterization of the relation between the metal content of the ionized gas with the ratios of fluxes emitted by different ions, should also cover all physical mechanisms other than and , including dust-to-gas mass ratio or gas geometry. In any case, the functional observed anticorrelation between O/H and log , that implicitly includes all these factors as they are not varied in the models, is still useful, given that, firstly, it still allows a precise determination of chemical abundances, reducing the number of explored models in the grid, and, secondly, it also allows a consistent comparison between objects in conditions where not all the mechanisms affecting the excitation can be independently fixed.
@@ 4.3 Correcting the [S ii] emission excess in MaNGA
One of the drawbacks that the S abundance derivation may have in samples of objects with very low spatial resolution comes from the confirmed fact that a non-negligible fraction of the collected [S ii] emission presents an excess in relation to what it would be expected just accounting for the gas ionized by a resolved young massive stellar clusters (e.g. Zurita et al. 2000; Haffner et al. 2009). This diffuse ionized gas (DIG) emission may have a non-negligible impact on the determination of chemical abundances (e.g. Kumari et al. 2019), and it has also been observed in the specific case of the sample of selected MaNGA pointings studied here by Zinchenko et al. (2021), who realized that the N/O abundance ratios derived using HCm were much lower when the code was fed with [S ii] emission lines. Since the code uses as source of information to provide a solution for N/O both the emission-line ratios N2O2 and N2S2, which are tracers of this abundance ratio Pérez-Montero & Contini (2009), this implies that the [S ii] emission is overestimated in relation to what would correspond to the N/O abundance as solely traced by N2O2. This was later confirmed by Pérez-Montero et al. (2023), who studied the MaNGA sample in the context of the softness diagram and showed that the hardness of the field of radiation derived for these selected H ii regions was higher than the maximum considered for massive stars in a 60% of the sample when the S2S3 ratio was used. Instead, this proportion was lowered to values compatible with other H ii regions observed using narrow long-slit spectroscopy, such as CHAOS Berg et al. (2015), when [N ii] was used replacing [S ii].
Therefore, in order to provide a more precise derivation of the S abundance in the MaNGA sample, it is fundamental to quantify and correct the [S ii] excess , as [S ii] lines are necessary as input for our code.
@@
Several recipes have been proposed to correct the emission from DIG in galaxy disks with star formation (e.g. Vale Asari et al. 2019), but a consistent criterion that can be applied to a large number of objects like those in MaNGA without a detailed study of the geometry or the source of the incident field of radiation in each region is not straightforward (Weber et al., 2019). As an alternative, even though DIG emission also contributes to the emission of other low-excitation lines such as [N ii] (e.g. Blanc et al. 2009) or H recombination lines, we can consider that the excess is much lower than in the case of [S ii] in our sample of star-forming dominated selected objects. In that case, in order to minimize and correct the DIG emission in [S ii], we can use [N ii] and the resulting N/O ratios from hcm, as the N2O2 parameter is relatively unaffected by DIG emission (Zhang et al., 2017).
@@
In this way, taking as correct or at least weakly affected by DIG emission, those N/O values derived by HCm when only the lines involved in the N2O2 parameter are used, we can use these abundance ratios and a linear fit of them with the N2S2 parameter as predicted by the photoionization models presented in Pérez-Montero (2014) to derive the [S ii] flux not affected by DIG emission. The functional form of this fit is:
@@
where N2S2 = log([N ii] 6584 Å/[S ii] 6717+31 Å). This relation has a Pearson statistical significance r = 0.992 and a p-value = 0.0. When models considering a S/O ratio different than the solar value are also examined, this function does not lead to [S ii] fluxes deviating more than a 3% in relation with the fluxes derived by this fit, what is much less than the measured excess in most of the sample.
@@
In Figure 7, we show the difference between the total S abundance calculated from HCm using the corrected and the measured [S ii] lines as a function of the total oxygen abundance, even for those objects for which the corrected value is larger than the observed one. As it can be seen, there is a slight correlation to find larger deviations for higher metallicities, and in the over-solar regime the S abundances can be in average 0.3 dex larger if the [S ii] excess is not considered, underlying the importance of this correction even for star-forming dominated objects.
@@ 4.4 The relation between metallicity and S/O
Figure 8 displays the relation between the sulphur-to-oxygen abundance ratio, log(S/O), and the oxygen abundance, 12 + log(O/H), as derived by HCm for the MaNGA sample of star-forming selected pointings. Each blue dot represents a pointing in which all required emission lines were measured with sufficient signal-to-noise. A grey density map highlights the statistical distribution of the points. A red dashed horizontal line marks the solar S/O value (log(S/O) , Asplund et al. 2009), and a red dot indicates the solar oxygen abundance (12 + log(O/H) ). The figure also includes comparison samples from Díaz & Zamora (2022), based on the direct method: classical disk H ii regions (DHR, orange circles) and high-ionization H ii galaxies (HIIG, green squares). We also include the sample of star-forming objects studied by Pérez-Díaz et al. (2024a) using HCm-IR and mid-IR emission-lines (red triangles).
In addition, since the low-metallicity regime (12+log(O/H) 7.6) is poorly populated, by the MaNGA sample and also very few H ii galaxies taken from Díaz & Zamora (2022), we also included the sample of XMPs studied by Zinchenko et al. (2024) (blue diamonds) to cover this range. In this regime, we observe some scatter, although considering the uncertainties in the estimations, most of them are compatible with the S/O solar ratio. although a number of individual points show enhanced log(S/O). These outliers may reflect local chemical enrichment by Type II supernovae or other stochastic effects in metal-poor galaxies. Alternative scenarios include IMF variations or contributions from very massive stars (e.g., Goswami et al., 2024).
In the range 7.6 12+log (O/H) 8.0, both samples from MaNGA and Pérez-Díaz et al. (2024a), present a non-negligible number of objects with S/O ratios slightly over solar. However, when considering all data, the majority of the objects show an almost solar S/O ratio. This behavior is consistent with the expectation that both oxygen and sulphur, being produced mainly by core-collapse supernovae, evolve in lockstep.
@@
As suggested by Díaz & Zamora (2022), we can minimize the dependence of total metallicity on its depletion onto dust, by representing its relation with S/O using total sulphur abundance. This is shown in Figure 9 for the same samples described above, along with the corresponding S/O value using again a red dashed line and the solar S/H value at 7.12 (Asplund et al., 2009).
We note that this plot is consistent with the previous one, as the sample from Zinchenko et al. (2024) is clearly differentiated from the others by covering the low-metallicity regime in S/H. Interestingly, we observe to major trends. In the range 5.9 12+log(S/H) 6.5, there is essentially no trend, observing that objects cluster around slightly sub-solar S/O values. On the other hand, although DHRs and IR data tend to cluster around the solar S/O ratio for 12+log(S/H) below 7.12, there is a clear trend to find increasing log(S/O) for increasing 12+log(S/H) at higher abundances. In addition, the MaNGA sample seems to present several behaviors. For over-solar S abundances, the same increasing trend is observed, while for sub-solar S abundances, there are two trends: one that, as in the previously described samples, has average values clustering around the solar S/O ratio, and another one showing the same increasing S/O values observed in the high-metallicity range.
Thus, our results indicating that the increase of log(S/O) with the total sulphur abundance is much clearer than when it is studied as a function of total oxygen abundance, is an evidence of the growing depletion of this element into dust grains for higher metallicity, underlining that total sulphur abundance can be a better tracer of the overall metal content in these objects, and giving important clues on the amount of oxygen retired from the gas-phase.
This interpretation is reinforced by the fact that the MaNGA data has been corrected for the [S ii] excess, which is particularly relevant at high metallicities. As shown in Figure 7, and discussed by Pérez-Montero et al. (2023), DIG can lead to overestimated [S ii] fluxes in IFS data, and thus to lead to artificially higher sulphur abundances. The application of a correction based on the model-based N2S2–N2O2 relation mitigates this effect, reducing potential overestimates in log(S/O) by up to 0.3 dex at solar and over-solar O/H. Without such correction, the observed increase in log(S/O) at high metallicity would likely be stronger, overestimating the depletion issue for oxygen.

### v2

@@ 4.1 Comparison with the control sample
In Figure 3 we show the comparison between the abundances calculated using the data in Díaz & Zamora (2022) with the results from HCm when all possible lines, including auroral lines, are used as input by the code. We also list in Table LABEL:disp the mean offsets and the standard deviations of these comparisons. As it can be seen, the agreement in both plots is quite good, without any neither discontinuity nor break over a wide range of metallicity, and only with some little deviations that can be attributed to several factors including the atomic coefficients, the assumed relations between the electron temperatures associated to each ion, or the assumed ICFs for each element. These factors are arbitrary and can lead to deviations that are usually lower than the observational uncertainties.
In absence of any auroral line, which is the most common situation in weak regions or for metal-rich objects, where the cooling is much more efficient and the emissivities of all CELs is lower, the code can also provide a solution with its corresponding uncertainty. However, additional assumptions must be taken in order to provide a solution consistent with the direct method. As discussed in Pérez-Montero (2014), this is done by assuming an empirical relation between O/H and to constrain the grids, in such a way that models with lower metallicity have in average higher excitation and, on the contrary, low excitation regions are compatible with metal-rich regions. This assumption is strictly empirical and its interpretation, as it will be discussed in next subsection is subject to uncertainties that we partially consider in our code (i.e. by adopting ranges of variation in the adopted empirical relation),but it helps breaking the degeneracy of certain emission-line ratios such as R23, with metallicity. A similar constrain is considered by the code when N/O cannot be calculated, so [N ii] emission-lines can only be used to derive O/H after assuming a certain arbitrary relation between O/H and N/O. As discussed in Pérez-Montero (2014) and several other works describing HCm, this assumption is usually made in all strong-line methods based on N ii emission lines in the optical, so it can lead to non-negligible mistakes if the studied object does not follow the expected adopted relation. In any case, in order to let the user to explore additional uncertainties owing to the adoption of different relations, these can be edited or replaced in the code to account for different families of objects or prescriptions. For instance, as studied in Pérez-Montero et al. (2021), for EELGs, the assumed relation between O/H and is slightly different, pointing towards slightly higher values for lower metallicity.
@@
4.2 An interpretation of the empirical relation between O/H and
@@
One of the foundations of the above described methodology using the HCm code to derive chemical abundances when no nebular-to-auroral line ratios are provided in the input is the assumed empirical anticorrelation between O/H and , what is used by the code constraining the grid of models to avoid the double-valued relations of certain emission-line ratios with O/H (e.g. R23, Pagel et al. 1979). Although not universal and highly dependent on the studied class of objects, this assumption comes from a widely observed relation between certain ratios of consecutive emission-line fluxes and the chemical content of the gas. In this subsection, we try to make a deeper examination on the physical basis of this assumption, taking advantage of the direct relation between certain emission-line ratios with the input parameters considered in our grids of photoionization models.
@@
In left panel of Figure 5 we show the relation between the values of O/H and found by HCm for the objects in the control sample for which at least one auroral line is used as an input. In this case, no previous assumption is required on the relation between O/H and to retrieve precise estimates of the chemical abundances, and thus no constrain is made to the used grid of models.
@@
As it can be seen, the code finds for the subsample from Díaz & Zamora (2022) used in this work the expected anticorrelation already found for a larger sample in Pérez-Montero (2014), in such a way that objects with a lower metallicity have in average higher values for log . This relation is represented by the solid line shown in the same figure and also encompasses The set of models used by the code to perform the calculations when no auroral line is given as input. In appearance the O/H- relation derived for our control sample is , somehow less clear than in the case of the sample studied in Pérez-Montero (2014) ( = -0.48), maybe owing to the higher weight in the sample analyzed here of the objects at large metallicity, for which the trend does not seem to exist, and that some objects lie out of the limits designed to constrain the grid of models. However, the agreement between the abundances derived by the code when no auroral lines are given using a constrained grid is quite reasonable, as can be seen in Fig. 4 and inspecting the offsets and dispersions listed in Table LABEL:disp.
@@
Focusing on this observed O/H - anticorrelation itself, there is not a consensus in the literature about if this correlation is real or has a physical basis. On one hand, several authors (e.g. Dopita & Evans 1986; Maier et al. 2006; Morisset et al. 2016; Garner et al. 2025) have reported the same anticorrelation in samples of H ii regions by fitting high-to-low excitation emission-line ratios with photoionization models. On the other hand, some other authors found a lack of correlation or even a positive correlation between O/H and (e.g. Kennicutt & Garnett 1996; Dors et al. 2011; Kreckel et al. 2019; Zinchenko et al. 2019) and/or alert on the strong dependence of the results on the initial assumptions made in the photoionization models used to derive these properties (e.g. Ji & Yan 2022).
@@
In the several papers describing the different versions of HCm, it is stated that the derived values are subject to the definition of the assumed geometry or the adopted input SED, so its final value is just a convention taken to compare heterogeneous samples or to fix an empirical relation between the metal content of the gas and the observed behaviour of emission-line ratios very dependent on the gas excitation, such as O2O3 or S2S3, as explained in the above sections. However, a grid of models exploring a variation of properties such as the shape of the incident SED, the geometry, or the assumed dust-to-gas mass ratio is required to correctly interpret from a physical basis the observed anticorrelation between metallicity and , as all these other properties have a non-negligible impact on the observed emission-line ratios.
@@
It has been demonstrated that making small modifications on the assumed incident SED of the models does not imply changes beyond the reported errors for the results from HCm on the derived abundances, what is corroborated by the agreement found when these results are compared with those coming from the direct method. For instance, as explored in Pérez-Montero (2014) variations of the age of the popstar Mollá et al. (2009) stellar atmospheres used in the models do not lead to deviations in the derived oxygen or nitrogen abundances larger than 0.02 dex. In addition, in Pérez-Montero et al. (2019a), for AGN, a change in the assumed parameter shaping the incident SED, do not alter the derived abundances. Finally, in Pérez-Montero et al. (2021), for EELGs, it is proven that changes in the properties of the assumed BPASS Eldridge et al. (2017) model atmospheres, such as age, IMF or upper mass limit, only imply variations in the derived log , but neither for N/O nor for O/H.
Nevertheless, a closer inspection of the real nature of the derived values in HCm can only be made by means of an alternative approach exploring the nature of the incident field of radiation. For this reason, in left panel of Figure 5 we use the denomination to denote the ionization parameter calculated by HCm to underline the fact that the used grid of models do not consider variations of the incident SED (i.e. they are calculated at a around fixed equivalent effective temperature, ) so the variation of certain emission-line ratios usually associated to variations of the excitation (e.g. O2O3, S2S3) are exclusively interpreted by the code as due to changes of according to this grid of models.
Alternatively, an analysis of considering variations of the incident field of radiation can be performed by using the code HCm-Teff Pérez-Montero et al. (2019b), with a very similar methodology to HCm, but using as inputs the metal content of the gas, and at least two emission-line ratios to simultaneously fix and , thus providing an additional perspective to interpret the observed anticorrelation between O/H and the obtained parameter when the SED is fixed. The study of the control sample studied in this work allows this study, as it has been selected to simultaneously have emission-line ratios, O2O3 and S2S3 involved in the softness diagram Pérez-Montero & Vílchez (2009), what allows HCm-Teff to extract both and .
@@
The anticorrelation between O/H and was already pointed by Pérez-Montero et al. (2019b) analyzing the H ii regions in NGC 628, M 51, and M 101, where the presence of a negative radial gradient of metallicity were obtained in coincidence with a positive radial gradient of the hardness of the incident field of radiation. This relation was also indirectly pointed out by other authors (e.g. Dopita et al. 2006; Eldridge et al. 2011), who state that the relation observed with when this is derived using emission-line ratios could be the consequence of a higher opacity in the atmospheres of the ionizing stars. At same time, other empirically found relations between metallicity and other properties of the ionizing stellar population have been found in other works, such as Initial Mass Function (IMF, Martín-Navarro et al. 2015), SFR or sSFR (e.g. Dopita et al. 2014; Mingozzi et al. 2020), or age Pellegrini et al. (2020), underlining the anticorrelation between metallicity and the hardness of the incident field of radiation, and not for the ionization parameter.
Nevertheless, a more precise and robust characterization of the relation between the metal content of the ionized gas with the ratios of fluxes emitted by different ions, should also cover all physical mechanisms other than and , including dust-to-gas mass ratio or gas geometry. In any case, the functional observed anticorrelation between O/H and log , that implicitly includes all these factors as they are not varied in the models, is still useful, given that, firstly, it allows a precise determination of chemical abundances, reducing the number of explored models in the grid, and, secondly, it also allows a consistent comparison between objects in conditions where not all the mechanisms affecting the excitation can be independently fixed.
@@ 4.3 Correcting the [S ii] emission excess in MaNGA
One of the drawbacks that the S abundance derivation may have in samples of objects with very low spatial resolution comes from the confirmed fact that a non-negligible fraction of the collected [S ii] emission presents an excess in relation to what it would be expected just accounting for the gas ionized by a resolved young massive stellar clusters (e.g. Zurita et al. 2000; Haffner et al. 2009). This diffuse ionized gas (DIG) emission may have a non-negligible impact on the determination of chemical abundances (e.g. Kumari et al. 2019), having a non-negligible weight in different galaxy surveys based on integral field spectroscopy like CALIFA (Lacerda et al., 2018) or MaNGA (Belfiore et al., 2016). Moreover, the impact of the excess in the [S ii] emission on the determination of abundances was already documented in the specific case of the sample of selected MaNGA pointings studied here by Zinchenko et al. (2021), who realized that the N/O abundance ratios derived using HCm were much lower when the code was fed with [S ii] emission lines, allowing that both N2O2 and N2S2 parameters were simultaneously used. Since the code uses these two emission-line ratios as source of information to provide a solution for N/O (Pérez-Montero & Contini, 2009), this implies that the [S ii] emission is overestimated in relation to what would correspond to the N/O abundance as solely traced by N2O2, relatively unaffected by DIG emission (Zhang et al., 2017).
@@
Several recipes have been proposed to correct the emission from DIG in galaxy disks with star formation (e.g. Vale Asari et al. 2019), but a consistent criterion that might be applied to a large number of objects like those in MaNGA without a detailed study of the geometry or the source of the incident field of radiation in each region is not straightforward (Weber et al., 2019). As an alternative, even though DIG emission also contributes to the emission of other low-excitation lines such as [N ii] (e.g. Blanc et al. 2009) or H recombination lines, we can consider that the excess due to the DIG in these lines is much lower than in the case of [S ii] in our sample of star-forming dominated selected objects. This is partially confirmed for the sample studied here by the results from Pérez-Montero et al. (2023), who studied the MaNGA sample in the context of the softness diagram and showed that the hardness of the field of radiation derived for these selected H ii regions is higher than the maximum considered for massive stars in a 60% of the sample when the S2S3 ratio is used. Instead, this proportion is lowered to values compatible with other H ii regions observed using narrow long-slit spectroscopy, such as CHAOS Berg et al. (2015), when [N ii] is used replacing [S ii].
@@
Therefore, in order to provide a more precise derivation of the S abundance in the MaNGA sample, it is fundamental to quantify and correct the [S ii] excess, as [S ii] lines are necessary as input for our code to derive S, while this line is not used to derive other abundances, like O/H or N/O, whose results are then much less affected by this contribution to other lines. In this way, assuming that the lines involved in the N2O2 parameter are much less affected by DIG emission, we can minimize and correct the DIG emission of [S ii], by taking the N/O values derived by HCm when only [N ii] and [O ii] are used. Later, these abundance ratios can be utilized to derive the fraction of [S ii] flux not affected by DIG by means of the relation of them with the N2S2 parameter. as predicted by the photoionization models presented in Pérez-Montero (2014). The functional form of this fit is:
@@
where N2S2 = log([N ii] 6584 Å/[S ii] 6717+31 Å). This relation has a Pearson statistical significance r = 0.992 and a p-value = 0.0. When models considering a S/O ratio different than the solar value are also examined, this function does not lead to [S ii] fluxes deviating more than a 3% in relation with the fluxes derived by this fit, what is much less than the measured excess in most of the sample. We then took for our analysis the resulting corrected [S ii] derived from this relation.
@@
In order to illustrate the impact of this correction on S abundances, we show in Figure 7, the difference between the total S abundances using the measured and the corrected [S ii] lines calculated from HCm. The difference is presented as a function of the total oxygen abundance, even for those objects for which the derived S/H abundance based on the corrected [S ii] line is larger than the value based on the measured one. As it can be seen, there is a slight correlation to find larger deviations for higher metallicities, and in the over-solar regime the S abundances can be in average 0.3 dex larger if the [S ii] excess is considered, underlying the importance of this correction even for star-forming dominated objects.
@@ 4.4 The relation between metallicity and S/O
Once we have a methodology to derive sulphur abundances that can be applied to a large and representative sample of objects, regardless of the fact that the direct method can be used, we examine the overall behaviour of the relation between metallicity and S/O ratio over a wide range of metallicity. Even though the MaNGA sample presents certain objects for which both [O iii] and [S iii] auroral lines could be measured with enough signal-to-noise ratio (around 1131, a 0.56% of the sample), we only considered strong nebular lines for the whole sample for our calculations, under the above discussed assumption of an empirical relation between O/H and log . Anyway, we checked that the derived abundances in this sample using the auroral lines did not deviate more than the values already discussed.
Figure 8 displays the relation between the sulphur-to-oxygen abundance ratio, log(S/O), and the oxygen abundance, 12 + log(O/H), as derived by HCm for the MaNGA sample of star-forming selected pointings. Each blue dot represents a pointing in which all required emission lines were measured with sufficient signal-to-noise. A grey density map highlights the statistical distribution of the points. while the black circles joined by a black line represents median values per bin in the sample. A red dashed horizontal line marks the solar S/O value (log(S/O) , Asplund et al. 2009), and a red dot indicates the solar oxygen abundance (12 + log(O/H) ).
The figure also includes comparison samples from Díaz & Zamora (2022), based on the direct method: classical disk H ii regions (DHR, orange circles) and high-ionization H ii galaxies (HIIG, green squares). We also include the sample of star-forming objects studied by Pérez-Díaz et al. (2024a) using HCm-IR and mid-IR emission-lines (red triangles). In addition, since the low-metallicity regime (12+log(O/H) 7.6) is poorly populated, by the MaNGA sample and also very few H ii galaxies taken from Díaz & Zamora (2022), we also included the sample of Extremely Metal Poor Galaxies (XMPs; O/H O/H⊙/10) studied by Zinchenko et al. (2024) (blue diamonds) to cover this range.
In this regime, we observe some scatter that, considering the uncertainties in the estimations, is compatible with the S/O solar ratio, although a number of individual points show enhanced log(S/O), a trend which is also traced in the MaNGA sample. These outliers may reflect local chemical enrichment by Type II supernovae or other stochastic effects in metal-poor galaxies. Alternative scenarios include IMF variations or contributions from very massive stars (e.g., Goswami et al., 2024). In the range 7.6 12+log (O/H) 8.0, both samples from MaNGA and Pérez-Díaz et al. (2024a), present a non-negligible number of objects with S/O ratios slightly over solar. However, when considering all data, the majority of the objects show an almost solar S/O ratio. This behavior is consistent with the expectation that both oxygen and sulphur, being produced mainly by core-collapse supernovae, evolve in lockstep.
@@
As suggested by Díaz & Zamora (2022), we can minimize the dependence of total metallicity on its depletion onto dust, by representing its relation with S/O using total sulphur abundance, which is thought to not to be affected by this. This is shown in Figure 9 for the same samples described above, along with the corresponding S/O value using again a red dashed line and the solar S/H value at 7.12 (Asplund et al., 2009).
We note that this plot is consistent with the previous one, as the sample from Zinchenko et al. (2024) is clearly differentiated from the others by covering the low-metallicity regime in S/H. Interestingly, we observe to major trends. In the range 5.9 12+log(S/H) 6.5, there is essentially no trend, observing that objects cluster around slightly sub-solar S/O values, and without any object with oversolar S/O values in this regime, what could point to that the objects simultaneously showing low O/H and high h S/O ratios could have very large dust contents. On the other hand, although DHRs and IR data tend to cluster around the solar S/O ratio for 12+log(S/H) below 7.12, there is a clear trend to find increasing log(S/O) for increasing 12+log(S/H) at higher abundances. In addition, the MaNGA sample seems to present several behaviors. For over-solar S abundances, the same increasing trend is observed, while for sub-solar S abundances, there are two trends: one that, as in the previously described samples, has average values clustering around the solar S/O ratio, and another one showing the same increasing S/O values observed in the high-metallicity range.
Thus, our results indicating that the increase of log(S/O) with the total sulphur abundance is much clearer than when it is studied as a function of total oxygen abundance, is an evidence of the growing depletion of this element into dust grains for higher metallicity, underlining that total sulphur abundance can be a better tracer of the overall metal content in these objects, and giving important clues on the amount of oxygen retired from the gas-phase. This interpretation is reinforced by the fact that the MaNGA data has been corrected for the [S ii] excess, which is particularly relevant at high metallicities. As shown in Figure 7, and discussed by Pérez-Montero et al. (2023), DIG can lead to overestimated [S ii] fluxes in IFS data, and thus to lead to artificially higher sulphur abundances. The application of a correction based on the model-based N2S2–N2O2 relation mitigates this effect, reducing potential overestimates in log(S/O) by up to 0.3 dex at solar and over-solar O/H. Without such correction, the observed increase in log(S/O) at high metallicity would likely be stronger, overestimating the depletion issue for oxygen.